A metal crystallises in two phases, fcc and bcc, whose unit cell lengths are 3.5 Armstrog and 3.0 Armstrong respectively. The ratio of density of fcc and bcc is?

Grade:12

3 Answers

Adarsh
6 years ago
Dear Sai,
the answer will be 1.259 .  Use the formula for density, = z.M/Na.a3. .
Thanks and Regards
Adarsh
Khimraj
6 years ago
Density is given by
d = (z*M)/(a3*NA)
given that a1 = 3.5 A0 and a2 = 3.0 A0
No of atom per unit cell of fcc z1 = 4
No of atom per unit cell of fcc z2 = 2
Sonratio of density of fcc and bcc
d1/d2 = (z1/z2)*(a2/a1)3 = (4/2)*(3/3.5)3 = 2*(6/7)3 =432/343.
Hope it clears. If you like answerthen please approve it.
Anjali kumari
5 years ago
As density= Z..M/Na.a^3. You will get 432:343.If you like the answer please approve it.Thank you. Gusto do TCP go gdp cpc